OpenStudy (anonymous):

What is 900 feet per day converted into yards per hour

OpenStudy (anonymous):

900 feet per day converted into yards per hour is: 12.5 yards per hour **900 ft/day x 1 yard/3ft x 1 day/ 24 hours = 12.5 yards/hour

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Np

OpenStudy (anonymous):

900 feet per day \[=\frac{ 900 feet }{ 1 day }=\frac{ \frac{ 900 }{ 3 } yards }{ 1*24hours }=\frac{ 300~ yards }{ 24~ hours }\] =? yards per hour.